# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v5

@@ 4.2.2 Variational Autoencoders (VAEs)
Early methods for uncertainty quantification in medical imaging primarily utilized 2D slices from three-dimensional (3D) datasets, leading to a potential loss of critical spatial information and subtle nuances often necessary for accurate delineation. This limitation has spurred research into 3D probabilistic segmentation techniques with ELBO-based models, aiming to preserve the integrity of entire 3D structures. Initial works (Chotzoglou & Kainz, 2019; Long et al., 2021a) demonstrate that the PU-Net can be adapted by replacing all 2D operations with their 3D variants. Crucially, the fusion of the latent sample with 3D extracted features is achieved through a 3D tiling operation. Viviers et al. (2023c) additionally augment a Normalizing Flow to the posterior density for improved expressiveness, as discussed in Section 4.2.2. Further enhancements to the architecture include the implementation of the 3D HPU-Net (Saha et al., 2020), an updated feature network incorporating the attention mechanisms, a nested decoder, and different reconstruction loss components tailored to specific applications (Saha et al., 2021a).
@@
It can be relevant to model the annotators independently in cases with consistent annotator-segmentation pairs in the dataset. This can be achieved by conditioning the learned densities on the annotator itself (Gao et al., 2022; Schmidt et al., 2023). For example, features of a U-Net can be combined with samples from annotator-specific Gaussian-distributed posteriors (Schmidt et al., 2023). Considering the approach from Gao et al. (2022), generating a segmentation mask is achieved by first sampling an annotator from a categorical prior distribution $\mathcal{C}(\pi_{k}(\mathbf{x}))$, governed by the image conditional parameters $\pi_{k}(\mathbf{x})$ for the $k$-th annotator. Then, samples are taken from its corresponding prior density as $\mathbf{z}_{k}\sim p_{k}(\mathbf{z}_{k})$ to reconstruct a segmentation through an image-conditional decoder $)\*y=F(\mathbf{x},\mathbf{z}_{k}$ ). The parameters $\pi_{k}(\mathbf{z}_{k})$ can also be used to weigh the corresponding predictions to express the uncertainty in the prediction ensemble. Additionally, consistency between the model and ground-truth distribution is enforced through an optimal transport loss between the set of predictions and labels.
@@ 4.2.3 Denoising Diffusion Probabilistic Models (DDPMs)
Figure 10: Illustration of the Diffusion Probabilistic Models (Ho et al., 2020). The model learns to remove noise that has been gradually added to the input image.
Recent developments in generative modeling have resulted in a family of models known as Denoising Diffusion Probabilistic Models (Ho et al., 2020; Song et al., 2022; 2020). Such models are especially renowned for their expressive power by being able to encapsulate large and diverse datasets. While several perspectives can be used to introduce the DDPMs, we build upon the earlier discussed HVAE (Section 4.2.2) to maintain cohesiveness with the overall manuscript. Specifically, we describe three additional modifications to the HVAE (Luo, 2022). Firstly, the latent dimensionality is set equal to the spatial data dimensions, i.e. $d\,$=$\,D$. As a consequence, redundant notation of $\mathbf{Z}$ is removed and $\mathbf{Y}$ is instead subscripted with $t\in\{1,...,T\}$, indicating the encoding depth, where $\mathbf{Y}_{0}$ is the initial segmentation mask. Secondly, the encoding process (or forward process) is predefined as a linear Gaussian model such that
@@
where
$q(\mathbf{y}_{t}|\mathbf{Y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\bm{\mu}=\sqrt{\alpha_{t}}\mathbf{Y}_{t-1},\bm{\Sigma}=(1-\alpha_{t})\cdot\mathbf{I}),$ (30)
with noise schedule $\bm{\alpha}=\{\alpha_{t}\}^{T}_{t=1}$. Then, the decoding or reverse process can be learned through $p_{\phi}(\mathbf{Y}_{t-1}|\mathbf{Y}_{t},\mathbf{x})$. The ELBO for this objective is defined as
@@
As denoted, the regularization term is assumed to be zero, since it is assumed that a sufficient amount of steps $T$ are taken such that $q(\mathbf{y}_{T}|\mathbf{y}_{0})$ is approximately normally distributed. With the reparameterization trick (Kingma & Welling, 2013), the forward process is governed by random variable $\bm{\epsilon}\sim\mathcal{N}(0,1)$. As such, the KL divergence in the second term can be interpreted as predicting $\mathbf{Y}_{0}$, the source noise $\bm{\epsilon}$ or the score $\nabla_{\mathbf{Y}_{t}}\log q(\mathbf{y}_{t})$ (gradient of the data log-likelihood) from $\mathbf{Y}_{t}$, depending on the parameterization. This is in almost all instances approximated with a U-Net (Ronneberger et al., 2015).
[…]
@@ 5 Parameter modeling
The key distinction between epistemic and aleatoric uncertainty lies in their origins: epistemic uncertainty arises from model ignorance (i.e. related to the parameters), whereas aleatoric uncertainty reflects inherent statistical variability in the data. Consequently, unlike aleatoric uncertainty, epistemic uncertainty should not only be quantified but ideally also minimized. Epistemic uncertainty can be further categorized into two distinct types (Hüllermeier & Waegeman, 2021). The first type pertains uncertainty related to the capacity of the model. For example, under-parameterized models or approximate model posteriors, that can become too stringent to appropriately resemble the true posterior. The ambiguity on the best parameters given the limited capacity induces uncertainty in the learning process, which is also known as model uncertainty. Nevertheless, given the complexity of contemporary parameter-intensive CNNs, the model uncertainty is often assumed to be negligible. A more significant contribution to the epistemic uncertainty is due to the limited data availability, known as approximation uncertainty, and can often be reduced by collecting more data.
As discussed in Section 3, epistemic uncertainty is captured by modeling the posterior distribution defined in Equation (1). However, evaluating the true Bayesian posterior is impeded by the intractability of the data likelihood in the denominator. Hence, extensive efforts have been taken to obtain viable approximations, such as using Mean-Field Variational Inference Blundell et al. (2015), Markov Chain Monte Carlo (MCMC) Neal (2012), Monte-Carlo Dropout Gal & Ghahramani (2016); Kingma et al. (2015), Model Ensembling Lakshminarayanan et al. (2017), Laplace approximations Mackay (1992), Stochastic Gradient MCMCs Korattikara Balan et al. (2015); Springenberg et al. (2016); Welling & Teh (2011), assumed density filtering Hernández-Lobato & Adams (2015) and expectation propagation Hasenclever et al. (2017); Louizos & Welling (2016). We refer to any neural network that approximates the Bayesian posterior over model parameters as a Bayesian Neural Network (BNN). Unlike modeling the output distribution directly, sampling parameters from the posterior naturally induces spatial coherence in the output. As a result, existing methods are often simple extensions of conventional classification BNNs, rather than leveraging techniques tailored for segmentation tasks. This section discusses three methods for modeling epistemic uncertainty. The most commonly used approach is Variational Inference (VI) over model parameters (Section 5.1). Test-time augmentation, often mistakenly regarded as an aleatoric method, is addressed in Section 5.3. Finally, Laplace Approximations are examined in Section 5.2.
@@ 5.1 Variational Inference
Consider a simpler, tractable density $q(\bm{\theta}|\bm{\eta})$, parameterized by $\bm{\eta}$ (e.g. a gaussian with $\bm{\eta}=\{\mu,\sigma\}$), to approximate posterior $p(\bm{\theta}|\mathbf{y},\mathbf{x})$. Then, we can employ Variational Inference (VI) w.r.t. to the parameters, by minimizing the Kullback-Leibler (KL) divergence between the true and approximated Bayesian posterior as
@@
where the parameter-independent terms are constant and therefore excluded from notation. In the case of a deterministic encoder, i.e. $q(\bm{\theta}|\bm{\eta})=\delta(\bm{\theta}-\bm{\theta}^{*})$, the formulation collapses to the MAP estimate in Equation (5). A popular choice for the approximated variational posterior is the Gaussian distribution, i.e. a mean $\mu$ and covariance $\sigma$ parameter for each element of the convolutional kernel, usually with zero-mean unitary Gaussian prior densities. However, the priors can be also learned through Empirical Bayes (Bishop, 1995). Furthermore, backpropagation is possible with the reparameterization trick (Kingma & Welling, 2013) and within this context, the procedure is referred to as Bayes by Backprop (BBB) (Blundell et al., 2015) and has later been improved with the Local Reparameterization trick (Kingma et al., 2015). In addition to approximating the posterior, sampling from it yields multiple parameter permutations, effectively enriching the model’s hypothesis space, a phenomenon also referred to as model combination (Minka, 2000; Clarke, 2003; Lakshminarayanan, 2016). In most cases, VI is not performed explicitly. Instead, simpler techniques are employed, with the resulting parameter permutations hypothesized to serve as a proxy. We discuss two such methods: Monte Carlo Dropout (Section 5.1.1) and Ensembling (Section 5.1.2). See Figure 11 for an illustration of each method. All of these methods have been widely applied to segmentation. However, unlike aleatoric uncertainty modeling, their implementations are often more generic and not specifically tailored to segmentation tasks. Therefore, we provide only a brief overview here, followed by a more in-depth discussion in later sections related to various applications.
@@
Figure 11: Three types of techniques are visualized to sample parameters of the convolutional kernels, (a) an approximation of the parameter density can be made, (b) taking samples can be mimicked with dropout during test time, or, (c) an ensemble of $N$ different configurations can be explicitly modeled.
@@ 5.1.1 Monte Carlo Dropout
Dropout is a common technique used to regularize neural networks (Wager et al., 2013) by randomly “switching off” nodes of the neural network. Using Dropout can also be interpreted as a first-order equivalent $L_{2}$ regularization with additionally transforming the input with the inverse diagonal Fisher information matrix (Wager et al., 2013). Furthermore, with Monte-Carlo Dropout (MC Dropout), the random node switching is continued during test time, effectively sampling new sets of parameters, mimicking samples from an implicit parameter distribution $q(\tilde{\bm{\theta}}|\bm{\theta},p)$, defined as
@@
with probability $p$ and vector $\mathbf{n}$ operating element-wise on the parameters. It has been shown that MC Dropout can be interpreted as approximate VI in a Deep Gaussian Process (Gal & Ghahramani, 2016). In this manner, such a method is able to provide multimodal estimates of the model uncertainty. As noted by Gal et al. (2017a), the variance in the model output is primarily governed by the magnitudes of the weights rather than the dropout rate $p$. The dropout rate is usually optimized through grid search or simply set to $p=0.5$. The relationship between $p$ and the magnitude of the model weights has also be exploited to probabilistically prune neural networks Gonzalez-Carabarin et al. (2022). Gal et al. (2017a) propose to additionally learn $p$ using gradient-based methods, known as Concrete Dropout, to increase the influence of $p$. As the name suggests, a continuous approximation to the discrete distribution is used, known as the Concrete distribution (Maddison et al., 2016; Jang et al., 2016), to enable path-wise derivatives through $p$. Also, a generalization of MC dropout has been proposed, where the weights of the network, rather that its hidden output, are set to zero (Mobiny et al., 2021).
@@ 5.2 Laplace approximation
with $\bm{\theta}_{\text{MAP}}$ being the Maximum a Posteriori solution and Hessian matrix $\mathbf{H}=-\nabla^{2}\log h(\bm{\theta})|_{\bm{\theta}_{\text{MAP}}}$, i.e. the second derivative of the loss function evaluated at $\bm{\theta}_{\text{MAP}}$. This approach, known as the Laplace Approximation (LA) (Laplace, 1774), has the merit that it can be performed post-hoc to .e.g. pretrained neural networks. However, estimating the Hessian can become infeasible, as computation scales quadratically with the model parameters. This is usually circumvented by treating neural network partly probabilistic and/or approximating the Hessian with a more simplified matrix structure (Martens & Grosse, 2015; Botev et al., 2017; Daxberger et al., 2021).
@@ 5.3 Test-time augmentation
An image $\mathbf{X}$ can be understood as a one-of-many visual representations of the object of interest. For example, systematic noise, translation or rotation result in many realistic variations that approximately retain image semantics. Hence, data augmentation (Krizhevsky et al., 2012) has been used at test time (explaining the name test-time augmentation, or TTA), argued to obtain uncertainty estimates by efficiently exploring the locality of the likelihood function (Ayhan & Berens, 2022). By randomly augmenting input images with invertible transformation $\tilde{\mathbf{x}}=T_{\zeta}(\mathbf{x})$, with transformation parameters $\zeta$, a prediction is obtained with $\tilde{\mathbf{y}}=f_{\bm{\theta}}(\tilde{\mathbf{x}})$ and can then be inverted through $\mathbf{y}=T_{\zeta}^{-1}(\tilde{\mathbf{y}})$. Repeatedly performing this procedure results in a set of segmentation masks, which serve as an estimate of $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$. In segmentation literature, it has been hypothesized that TTA encapsualtes the aleatoric uncertainty in a better way (Wang et al., 2019a; Ayhan & Berens, 2022; Zhang et al., 2022a; Wang et al., 2019b; Rakic et al., 2024; Whitbread & Jenkinson, 2022; Roshanzamir et al., 2023). However, literature opposing this also exist and even suggests that TTA enables the modeling of epistemic uncertainty (Hu et al., 2022). Kahl et al. (2024) draw parallels with BNNs and demonstrate this through the (by now) familiar decomposition of the predictive uncertainty
where $t$ are sampled augmentations from the input space $\mathcal{T}$. Aleatoric uncertainty is captured because a perfectly trained model produces consistent predictions across augmentations of known data. Epistemic uncertainty arises when the mutual information between the predicted label and augmentation is greater than zero for unseen data, indicating the model has not yet learned invariance for such inputs.
6 Applications
This section briefly provides literature that incorporates uncertainty into downstream tasks using segmentation models. These tasks can be broadly categorized into four main groups: estimating the segmentation mask distribution subject to observer variability (Section 6.1), reducing labeling costs using Active Learning (Section 6.2), model introspection (i.e. ability to self-assess Section 6.3) and improved generalization (Section 6.4). For quick reference, Table 5 summarizes the types of dataset domains (e.g., medical, outdoor scenes) associated with various downstream tasks.

### v8

@@ 4.2.2 Variational Autoencoders (VAEs)
Early methods for uncertainty quantification in medical imaging primarily utilized 2D slices from three-dimensional (3D) datasets, leading to a potential loss of critical spatial information and subtle nuances often necessary for accurate delineation. This limitation has spurred research into 3D probabilistic segmentation techniques with ELBO-based models, aiming to preserve the integrity of entire 3D structures. Initial works (Chotzoglou and Kainz, 2019; Long et al., 2021b) demonstrate that the PU-Net can be adapted by replacing all 2D operations with their 3D variants. Crucially, the fusion of the latent sample with 3D extracted features is achieved through a 3D tiling operation. Viviers et al. (2023a) additionally augment a Normalizing Flow to the posterior density for improved expressiveness. Further enhancements to the architecture include the implementation of the 3D HPU-Net (Saha et al., 2020), an updated feature network incorporating the attention mechanisms, a nested decoder, and different reconstruction loss components tailored to specific applications (Saha et al., 2021a).
@@
It can be relevant to model the annotators independently in cases with consistent annotator-segmentation pairs in the dataset. This can be achieved by conditioning the learned densities on the annotator itself (Gao et al., 2022; Schmidt et al., 2023). For example, features of a U-Net can be combined with samples from annotator-specific gaussian-distributed posteriors (Schmidt et al., 2023). Considering the approach from Gao et al. (2022), generating a segmentation mask is achieved by first sampling an annotator from a categorical prior distribution $\mathcal{C}(\pi_{k}(\mathbf{x}))$, governed by the image conditional parameters $\pi_{k}(\mathbf{x})$ for the $k$-th annotator. Then, samples are taken from its corresponding prior density as $\mathbf{z}_{k}\sim p_{k}(\mathbf{z}_{k})$ to reconstruct a segmentation through an image-conditional decoder $y=F(\mathbf{x},\mathbf{z}_{k})$. The parameters $\pi_{k}(\mathbf{z}_{k})$ can also be used to weigh the corresponding predictions to express the uncertainty in the prediction ensemble. Additionally, consistency between the model and ground-truth distribution is enforced through an optimal transport loss between the set of predictions and labels.
@@ 4.2.3 Denoising Diffusion Probabilistic Models (DDPMs)
Recent developments in generative modeling have resulted in a family of models known as Denoising Diffusion Probabilistic Models (Ho et al., 2020; Song et al., 2022; 2020). Such models are especially renowned for their expressive power by being able to encapsulate large and diverse datasets. While several perspectives can be used to introduce the DDPMs, we build upon the earlier discussed HVAE (Section 4.2.2) to maintain cohesiveness with the overall manuscript. Specifically, we describe three additional modifications to the HVAE (Luo, 2022). Firstly, the latent dimensionality is set equal to the spatial data dimensions, i.e. dd\,=D\,D. As a consequence, redundant notation of $\mathbf{Z}$ is removed and $\mathbf{Y}$ is instead subscripted with $t\in\{1,...,T\}$, indicating the encoding depth, where $\mathbf{Y}_{0}$ is the initial segmentation mask. Secondly, the encoding process (or forward process) is predefined as a linear Gaussian model such that
@@
with $q(\mathbf{y}_{t}|\mathbf{Y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\bm{\mu}=\sqrt{\alpha_{t}}\,\mathbf{Y}_{t-1},\bm{\Sigma}=(1-\alpha_{t})\cdot\mathbf{I}),$ and noise schedule $\bm{\alpha}=\{\alpha_{t}\}^{T}_{t=1}$. Then, the decoding or reverse process can be learned through $p_{\phi}(\mathbf{Y}_{t-1}|\mathbf{Y}_{t},\mathbf{x})$. The ELBO for this objective is defined as
@@
As denoted, the regularization term is ignored, since it is assumed that a sufficient amount of steps $T$ are taken such that $q(\mathbf{y}_{T}|\mathbf{y}_{0})$ is approximately normally distributed. With the reparameterization trick (Kingma and Welling, 2013), the forward process is governed by random variable $\bm{\epsilon}\sim\mathcal{N}(0,1)$. As such, the KL divergence in the second term can be interpreted as predicting $\mathbf{Y}_{0}$, the source noise $\bm{\epsilon}$ or the score $\nabla_{\mathbf{Y}_{t}}\log q(\mathbf{y}_{t})$ (gradient of the data log-likelihood) from $\mathbf{Y}_{t}$, depending on the parameterization. This is in almost all instances approximated with U-Net variants (Ronneberger et al., 2015; Song et al., 2020).
The proposed methodologies for segmentation vary in the conditioning of the reverse process on the input image. For instance, Wolleb et al. (2021) concatenate the input image with the noised segmentation mask. Wu et al. (2023b) insert encoded image features to the U-Net bottleneck. Additionally, information on predictions $Y_{t}$ at a time step $t$ is provided in the intermediate layers of the conditioning encoder. This is performed by applying the Fast Fourier Transform (FFT) on the U-Net encoding, followed by a learnable attention map and the inverse FFT. The procedure of applying attention on the spectral domain of the U-Net encoding has also been implemented with transformers in follow-up work (Wu et al., 2023a). Amit et al. (2021) also encode both current time step and input image, but combine the extracted features by simple summation prior to inferring it through the U-Net. It has also been proposed to model Bernoulli instead of Gaussian noise (Chen et al., 2023a; Zbinden et al., 2023; Rahman et al., 2023; Bogensperger et al., 2023).
Figure 10: Illustration of conditional Diffusion Probabilistic Models (Ho et al., 2020). The model learns to remove noise that has been gradually added to the input image.
Key takeaways: There is a large focus on improving amortized inference in VAE-based models (richer posteriors, hierarchies, refined objectives). The GAN-based CAR model has received limited follow-up. More contemporary work incline towards using DDPMs, where innovations largely center on input-conditioning strategies and employing categorical distributions in the noising process.
@@ 5 Parameter modeling
A significant advantage of this parameter-sampling approach, when compared to modeling the output distribution directly, is that it naturally induces spatial coherence across the output segmentations. As a result, existing methods are often simple extensions of uncertainty modeling in conventional classification, rather than leveraging techniques tailored for segmentation tasks. As established in Section 3, the posterior distribution from Equation (1) is used to generate a predictive distribution via Equation (2), suggesting parameter distribution modeling as an alternative to the modeling in feature space. We refer to any neural network that approximates the Bayesian posterior over model parameters as a Bayesian Neural Network (BNN). However, evaluating the true Bayesian posterior is impeded by the intractability of the data likelihood in the denominator. This section discusses the methods used in image segmentation for approximating the parameter distribution. The most commonly used approach is Variational Inference (VI) over model parameters (Section 5.1). Laplace Approximations are examined in Section 5.2. Test-Time Augmentation (TTA) is covered in Section 5.3 (we justify this non-trivial choice later in Section 7.1).
@@ 5.1 Variational Inference
Consider a simpler, tractable density $q(\bm{\theta}|\bm{\eta})$, parameterized by $\bm{\eta}$, to approximate posterior $p(\bm{\theta}|\mathbf{y},\mathbf{x})$. Variational Inference (VI) w.r.t. to the parameters can be employed, by minimizing the Kullback-Leibler (KL) divergence between the true and approximated Bayesian posterior as
@@
where the parameter-independent terms are constant and therefore excluded from notation. In the case of a deterministic encoder, i.e. $q(\bm{\theta}|\bm{\eta})=\delta(\bm{\theta}-\bm{\theta}^{*})$, the formulation collapses to the MAP estimate in Equation (5). A popular choice for the approximated variational posterior is the gaussian distribution, i.e., a mean $\mu$ and covariance $\sigma$ parameter for each element of the convolutional kernel, usually with zero-mean unitary gaussian prior densities. However, the priors can be also learned through Empirical Bayes (Bishop, 1995). Furthermore, backpropagation is possible with the reparameterization trick (Kingma and Welling, 2013) and within this context, the procedure is referred to as Bayes by Backprop (BBB) (Blundell et al., 2015) and has later been improved with the Local Reparameterization trick (Kingma et al., 2015). In addition to approximating the posterior, sampling from it yields multiple parameter permutations, effectively enriching the model’s hypothesis space, a phenomenon also referred to as model combination (Minka, 2000; Clarke, 2003; Lakshminarayanan, 2016). In most cases, VI is not performed explicitly. Instead, simpler techniques are employed, with the resulting parameter permutations hypothesized to serve as a proxy. We discuss two such methods: Monte Carlo Dropout (Section 5.1.1) and Ensembling (Section 5.1.2). See Figure 11 for an illustration how each method approximates the convolutional kernel distribution. All of these methods have been widely applied to segmentation. However, unlike feature-level distribution modeling, their implementations are often more generic and not specifically tailored to segmentation tasks. Therefore, we provide only a brief overview here, followed by a more in-depth discussion in later sections related to various applications.
@@
Figure 11: Three techniques for sampling convolutional kernels, with grayscale values indicating kernel weight magnitudes. These include: (a) Sampling an gaussian approximation parameterized by a mean and standard deviation kernel. (b) Applying dropout at inference time. (c) Training NN independently trained models, providing a discrete set of parameter configurations.
@@ 5.1.1 Monte Carlo Dropout
Dropout is a common technique used to regularize neural networks by randomly “switching off” nodes of the neural network. Using Dropout can also be interpreted as a first-order equivalent $L_{2}$ regularization with additionally transforming the input with the inverse diagonal Fisher information matrix (Wager et al., 2013). Furthermore, with Monte-Carlo Dropout (MC Dropout), the random node switching is continued during test time, effectively sampling new sets of parameters, mimicking samples from an implicit parameter distribution $q(\tilde{\bm{\theta}}|\bm{\theta},p)$, defined as
@@
with probability $p$ and vector $\mathbf{n}$ operating element-wise on the parameters. It has been shown that MC Dropout can be interpreted as approximate VI in a Deep Gaussian Process (Gal and Ghahramani, 2016). In this manner, such a method is able to provide multimodal estimates of the uncertainty. The dropout rate is usually optimized through grid search or simply set to $p=0.5$. The relationship between $p$ and the magnitude of the model weights can also be exploited to probabilistically prune neural networks (Gonzalez-Carabarin et al., 2022). As noted by Gal et al. (2017a), the variance in the model output is primarily governed by the magnitudes of the weights rather than the dropout rate $p$. Hence, they propose to additionally learn $p$ using gradient-based methods, known as Concrete Dropout, to increase the influence of $p$. As the name suggests, a continuous approximation to the discrete distribution is used, known as the Concrete distribution (Maddison et al., 2016; Jang et al., 2016), to enable path-wise derivatives through $p$. Also, a generalization of MC dropout has been proposed, where the weights of the network, rather than its hidden output, are set to zero (Mobiny et al., 2021).
@@ 5.2 Laplace approximation
with $\bm{\theta}_{\text{MAP}}$ being the Maximum a Posteriori solution and Hessian matrix $\mathbf{H}=-\nabla^{2}\log h(\bm{\theta})|_{\bm{\theta}_{\text{MAP}}}$, i.e. the second derivative of the loss function evaluated at $\bm{\theta}_{\text{MAP}}$. This approach, known as the Laplace Approximation (LA) (Laplace, 1774), has the merit that it can be performed post-hoc on pretrained neural networks. However, estimating the Hessian can become infeasible, as computation scales quadratically with the model parameters. This is usually circumvented by treating neural networks partly probabilistic and/or approximating the Hessian with a more simplified matrix structure (Martens and Grosse, 2015; Botev et al., 2017; Daxberger et al., 2021).
@@ 5.3 Test-time augmentation
[…]
Key takeaways: Compared with earlier methods, parameter distribution modeling is largely model agnostic and rarely incorporates segmentation-specific architectures. In practice, many implementations reduce to approximate Variational Inference techniques, such as MC Dropout or Ensembling, followed by methods like Test-Time Augmentation and Laplace Approximations. Echoing their origins in classical classification, these concepts apply broadly to other domains.
6 Tasks
This section reviews literature on the downstream utilization of uncertainty in segmentation models, classifying the corresponding tasks into four main groups. Furthermore, Table 5 in Section 6.5 explicitly connects the downstream tasks to the specific application domains where they have been successfully employed.
Tasks: Estimating Observer Variability, the range of plausible segmentation masks (Section 6.1) Reducing labeling costs using Active Learning (Section 6.2) Giving the model the ability to self-assess with Model Introspection (Section 6.3) Improving Model Generalization (Section 6.4)